North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size

Q6 - 29/30 June 2021

However, Tuesday was a warm sunny day for Bill Dobson, Nigel Hall, Ian McCall, Jon Bradley, Trevor Wilford and Ian Pearson. The first job of the day was to move Dame Vera Lynn, No.3672, onto No 6 Road in Deviation Shed, with Andrew Jeffery driving the 08 shunter. Before that could be done, Nos 5 and 80135 had to be moved north about a yard to allow enough room to get 3672 into the Shed. This also meant removing all our wooden packing from the back of No 6 Road and then re-stacking it again at the back of No 6 Road after the shunt.

On the Q6 itself, all the new axle box oil feed boxes have been fitted bar one on the leading righthand side which needs some attention. The outside of the cab and also inside the cab were sanded and rubbed down. The wooden seating was fitted and secured inside the cab. A start was made on rubbing down the window frames. Peter Whitaker called in for a discussion with Nigel Hall about fitting BR transfers onto the tender, a job with which Peter will assist Nigel in about two weeks' time after his holidays . The righthand side crosshead cotter was removed after a lot of persuasion with a bar and a big hammer, after being jammed since last week. Then the piston rod was parted from the crosshead using wedges and small pieces of packing, being slightly out of line. The rod was realigned with the crosshead and the cotter refitted. A small securing cotter was to be made and fitted on Wednesday.

On Wednesday, the team was Bill Dobson, Nigel Bill, Steve Hyman, Ed Bolam, Nigel Hall, Jon Bradley, Ian Pearson and, lastly, myself after calling in at Pickering on the way in. All the roller doors in Deviation Shed were serviced by NESS of Scarborough during the day.

More prep work continued on the outside of the cab in readiness for a coat of primer, while the black area inside the cab was given a coat of gloss. All the cab window frames were rubbed down and given a coat of varnish: one of the windows is away having new glass fitted. The new oil box on the right leading side is still not finished as the thread into the axle box needs retapping. The securing cotter for the right-hand crosshead cotter was made and fitted. On the left-hand side crosshead, the piston rod was lined up to allow the cotter pin to be fitted but, after hammering it in, it got stuck and had to be removed. So, after moving the loco into a convenient position to allow access through a spoke in the wheel, a heavy bar was used and a big hammer to knock the cotter out. After removal of the cotter, it was dressed up and re-fitted successfully. Now problems really started. When the bump test in the backwards direction started, the piston would not go all the way back, leaving about 1/2" of travel still to go. Nigel Hall had acquired an endoscope for another job, so the left-hand rear drain cock was removed and the camera was inserted up the orifice. Everyone had a look around on the endoscope screen to see if there were any foreign bodies between the back of the piston and the rear of the cylinder but nothing could be found . Further investigation will be necessary next week to try and resolve this problem.